# The Relative Contributions of Accretion Disk versus Jet to the Optical and Mid-infrared Variability of Seyfert Galaxies

### v1

@@ 3.3 Correlation between optical and mid-infrared variability amplitudes
It is expected that IR emission is reprocessed from optical/UV emission (Lu et al., 2016), meaning that the variability characteristics in the optical bands will be imprinted in the IR wavelength. To investigate such a scenario in the current sample, we selected Seyfert galaxies that exhibit optical variability and have a redshift-corrected variability amplitude in the MIR. For the r-band light curve, the criteria are satisfied by 14 gNLS1s, 8 ngNLS1s, and 9 gBLS1s in both the W1 and W2 bands. However, for the W2 band, only 6 ngNLS1s meet the criteria, while 8 ngNLS1s meet the criteria for the W1 band. Using this sample, we constructed correlation plots between and both and . A representative plot showing the correlated variability amplitudes of the r-band optical light curve with the MIR W1 and W2 bands for gNLS1s (left), ngNLS1s (middle), and gBLS1s (right) is presented in Fig. 4. To examine the relationship between MIR and optical variability amplitudes, we employed orthogonal distance regression (ODR) fitting on the data points in the variability amplitude correlation plot, accounting for uncertainties on both axes. Furthermore, we calculated the Pearson rank correlation coefficient () to quantify the correlation between the variability amplitudes for the gNLS1s, ngNLS1s, and gBLS1s samples across both the optical and MIR wavelengths. In this plot, the upper panels illustrate the correlation between and , while the lower panels show the correlation between and . For the r-band light curve, the results of the correlation analysis, such as slope, intercept, Pearson r for the W1 and W2 bands, are displayed in the upper left corner of each panel in Fig. 4.
@@
The number of sources used to estimate the statistics are tabulated inside the bracket ‘[]’. |
@@ 3.4 Color variability
The optical emission observed in AGNs arises from two primary components: the quasi-thermal emission originating from the accretion disk and the non-thermal synchrotron emission from the relativistic jet. Analyzing the color variability of the current sample can provide an effective way to differentiate these components that contribute to the overall flux. Similarly, IR emission, which is considered reprocessed optical/UV radiation (see, Storchi-Bergmann et al., 1992; Lu et al., 2016), offers a unique opportunity to examine the interplay of thermal and non-thermal processes in AGNs. However, to ensure accurate spectral analysis, simultaneous measurements in different bands must be obtained within a 30-minute interval, as described in Sect. 2. This criterion helps mitigate any uncertainties caused by temporal variations and ensures a reliable study of the spectral behavior. Consequently, the color variability analysis in both optical and MIR wavelengths is limited to a subset of sources that meet this stringent timing requirement. The details of the final samples satisfying these conditions are provided at the end of Sect. 2.
@@
For the samples of gNLS1s, ngNLS1s, and gBLS1s mentioned in the last paragraph of Sect. 2, we generated color-magnitude plots, including g-r vs. r, r-i vs. i, and W1-W2 vs. W2, using the available quasi-simultaneous data across different optical and MIR bands. Here, the magnitudes of the longer wavelength are chosen on the X-axis, and the difference between the magnitudes of shorter and longer wavelengths is considered for the color to make the color-magnitude analysis consistent across optical and MIR wavelengths. Representative examples of these long-term best-fit color-magnitude plots for the gNLS1-J130522.74511640.2 (left), ngNLS1-J144019.82502201.7 (middle), and gBLS1-J110249.85525012.7 (right) are presented in Fig. 5, corresponding to g-r vs. r, r-i vs. i, and W1-W2 vs. W2, respectively, for top, middle, and bottom. To examine the relationship between color and magnitude, we employed ODR fitting on the data points in the color-magnitude plots, accounting for uncertainties in both axes. Furthermore, we calculated to quantify the correlation between color and magnitude for the gNLS1s, ngNLS1s, and gBLS1s samples across both optical and MIR wavelengths.
@@
To qualitatively identify bluer-when-brighter (BWB) or redder-when-brighter (RWB) trends from color-magnitude diagrams, we classified sources based on values, considering indicative of a BWB trend and indicative of an RWB trend. For sources with values of between and 0.5, we assumed that there is no significant correlation (termed as NOT) between color and magnitude. The number of sources that exhibit trends of RWB, BWB, or NOT, along with their fractions relative to the sample size, is summarized in Table 3. The slope, intercept, and color trends of the color-magnitude diagrams g-r vs. r, r-i vs. i, and W1-W2 vs. W2 for the individual source are tabulated in Table 4.
@@
gNLS1 [17] | 6 (35.29%) | 1 (5.88%) | 10 (58.82%) | gNLS1 [10] | 5 (50.00%) | 0 (0.00%) | 5 (50.00%) | gNLS1 [22] | 15 (68.18%) | 0 (00.00%) | 7 (31.82%) |
ngNLS1 [128] | 78 (60.94%) | 0 (0.00%) | 50 (39.06%) | ngNLS1 [74] | 55 (74.32%) | 0 (0.00%) | 19 (25.68%) | ngNLS1 [176] | 169 (96.02%) | 0 (00.00%) | 7 (3.98%) |
gBLS1 [10] | 8 (80.00%) | 0 (0.00%) | 2 (20.00%) | gBLS1 [6] | 3 (50.00%) | 0 (0.00%) | 3 (50.00%) | gBLS1 [10] | 3 (30.00%) | 1 (10.00%) | 6 (60.00%) |
The number of sources used to estimate the statistics are tabulated inside the bracket ‘[]’. |
@@ 4 Results and discussion
In this study, we systematically investigated flux and color variability for redshift-matched samples of gNLS1s, ngNLS1s, and gBLS1s in optical and mid-infrared wavelengths. For this purpose, we utilize high-cadence optical light curves in the g, r, and i bands from ZTF and mid-infrared light curves in the W1 and W2 bands from WISE. To analyze flux variability in the optical wavelength, we initially applied -test using Eq. 1 to the g, r, and i band light curves to determine the presence or absence of intrinsic variability. Subsequently, was calculated exclusively for the variable sources using equation 2, whereas F was calculated for the entire sample, including both variable and non-variable targets, using equation 3. The mean and median values of and are tabulated in Table 1.
@@
From Table 1 and Fig. 2, it is evident that gBLS1s exhibit greater variability, as indicated by both and , compared to gNLS1s across the g, r, and i bands, whether considering mean or median values. This finding is consistent with previous studies on smaller and larger samples of gNLS1s and gBLS1s (Klimek et al., 2004; Ai et al., 2010, 2013; Rakshit & Stalin, 2017). As a class, NLS1s demonstrate lower optical variability amplitudes than BLS1s. Such a lower variability in NLS1s can be attributed to their higher Eddington ratios (REdd). Since NLS1s accrete at rates significantly higher than BLS1s (Boroson & Green, 1992; Peterson et al., 2000; Ojha et al., 2020), their optical emission is expected to have a relatively stronger thermal component from the accretion disk compared to the synchrotron emission (e.g., see Zhou et al., 2007; Paliya et al., 2014). However, the optical emission from AGNs, dominated by thermal processes, is less variable than the Doppler-boosted synchrotron emission from jets, as a result, the higher contribution of thermal emission in NLS1 due to its higher accretion rate can effectively suppress the amplitude of their optical variability (e.g., see Ojha et al., 2019). However, in contrast to gBLS1s and gNLS1s, ngNLS1s are scarcely variable across the g, r, and i bands with a mean DC () of 5% (see Table 1, see, e.g., Fig. 2), suggest their variability due to instabilities in the accretion disc. Furthermore, the stark contrast in variability characteristics between the gNLS1 and ngNLS1 samples reflected in their mean duty cycles of approximately 72% and 5%, respectively, can be attributed to the presence of relativistic jets in gNLS1s. These jets, which are responsible for the observed gamma-ray emission in gNLS1s, likely contribute significantly to the enhanced and more frequent variability exhibited by this subclass. The relativistic beaming effect associated with the jets amplifies both the emission and its variability, resulting in a markedly higher detection rate of flux variations compared to jet-absent ngNLS1s.
For the optical wavelength | For the mid-infrared wavelength | ||||||||||||
(g-r) color vs. magnitude | (r-i) color vs. magnitude | (W1-W2) color vs. magnitude | |||||||||||
SDSS Name | Galaxy | Slope | Intercept | Color | Slope | Intercept | Color | Slope | Intercept | Color | |||
type | trends | trends | trends | ||||||||||
J112758.87362028.4 | gNLS1 | 0.230.10 | 4.821.88 | 0.51 | RWB | 0.200.20 | 04.263.77 | 0.56 | RWB | 0.230.04 | 04.130.49 | 0.50 | RWB |
J130522.74511640.3 | gNLS1 | 0.520.05 | 9.040.87 | 0.55 | RWB | 0.400.10 | 06.771.65 | 0.54 | RWB | 1.090.07 | 14.350.80 | 0.65 | RWB |
J144318.56472556.7 | gNLS1 | 0.300.07 | 5.441.34 | 0.27 | NOT | 0.530.17 | 09.563.00 | 0.37 | NOT | 1.160.05 | 15.400.61 | 0.75 | RWB |
J000545.32063945.4 | ngNLS1 | 0.360.15 | 6.892.86 | 0.39 | NOT | 0.570.21 | 10.743.90 | 0.62 | RWB | 0.990.06 | 14.650.86 | 0.69 | RWB |
J000935.91295425.8 | ngNLS1 | 0.390.11 | 8.052.17 | 0.53 | RWB | 1.290.49 | 25.959.72 | 0.52 | RWB | 1.190.05 | 18.110.73 | 0.86 | RWB |
J002151.43322520.0 | ngNLS1 | 0.110.12 | 2.472.21 | 0.15 | NOT | 0.690.24 | 13.054.60 | 0.75 | RWB | 1.330.20 | 18.602.73 | 0.72 | RWB |
J110249.85525012.7 | gBLS1 | 0.310.05 | 6.270.91 | 0.56 | RWB | 0.260.12 | 05.402.27 | 0.68 | RWB | 0.450.03 | 07.230.49 | 0.63 | RWB |
J120951.76181006.9 | gBLS1 | 0.270.09 | 5.331.65 | 0.54 | RWB | 0.090.09 | 01.721.70 | 0.53 | RWB | 0.020.01 | 01.270.16 | 0.31 | NOT |
J164258.81394837.0 | gBLS1 | 0.170.01 | 03.230.18 | 0.79 | RWB | 0.030.02 | 0.150.38 | 0.20 | NOT | 0.270.06 | 04.070.68 | 0.51 | RWB |
@@
Notes: A portion of this table is presented here to display its form and content, however, the entire table is available in the online electronic version. |
@@
In contrast to the accretion disc based variability in ngNLS1s, the origin of resulted flux variabilities of gNLS1s and gBLS1s could either be due to accretion disc instabilities (e.g., Wiita et al., 1991; Chakrabarti & Wiita, 1993; Mangalam & Wiita, 1993) or fluctuations emerging within the jet (e.g., Marscher & Gear, 1985; Wagner & Witzel, 1995; Marscher, 2014). Therefore, we have adopted the variability amplitude cut of 0.4 mags suggested by Bauer et al. (2009) to distinguish between jet-based variability and accretion disc instability-based variability. Thus, we first checked for the variable targets in the samples of gNLS1s and gBLS1s, and found that 15 out of 17 variable gNLS1s (88%) and 10 out of 10 variable gBLS1s (100%) exhibit 0.4 mags for the r-band light-curves. For the g-/ i-band light-curves, such statistics are (89/75%) and (100/100%) for the samples of gNLS1s and gBLS1s, respectively. From this analysis, it appears that gBLS1s tend to be more variable compared to gNLS1s, which could be due to the larger observed contribution of jets more closely aligned towards the observer line of sight in gBLS1s (e.g. Safna et al., 2020). A comparatively small percentage of the variable targets with 0.4 mags in the sample of gNLS1s could also be due to a difference in their accretion rates as discussed above (see 2nd para). On the other hand, a similar kind of variability nature resulted from the samples of gNLS1s, ngNLS1s and gBLS1s in the mid-infrared W1 and W2 bands (see Table 1, Fig. 3), supporting a reprocessing of optical/UV emission from the central engine of Seyfert galaxies by the dusty torus to MIR wavelength. Furthermore, to strengthen the reprocessing scenario found here, we have checked the correlation of with and for optically variable targets having 0.1 in the mid-infrared wavelength for the samples of gNLS1s, ngNLS1s, and gBLS1s. This has resulted in a very good correlations ( 0.5) of with and for the r-band light curves in the samples of gNLS1s and gBLS1s both for W1 and W2 bands (see Fig. 4). However, such correlations resulted in a less significant equal to 0.46 and 0.43 for W1 and W2 bands for the correlation between and only in the sample of gNLS1s for the g-band light curves. Overall, for the g-band light curves, we have found stronger with mean for W1 and W2 bands () equal to 0.71 & 0.73 for the correlation of with and , respectively in gBLS1s compared to value equal to 0.45 & 0.63 of gNLS1s. Such statistics of found to be 0.70 & 0.87 for gBLS1s and 0.56 & 0.57 for gNLS1s when considering r-band light curves. Thus, it appears that values are consistently found to be stronger in the sample of gBLS1s as compared to the sample of gNLS1s for the correlation of with and . A comparatively less significant found in the sample of gNLS1s could be due to the higher accretion rate of gNLS1s which effectively suppresses the amplitude of variability. Contrary to the significant correlations found for the g-band and r-band light curves, such correlations are found to be weaker for the i-band light curves. On the other hand, no such correlation has been found in the sample of ngNLS1s, except for i-band light curves. Since i-band data are very scarcely sampled compared to g and r-band data, therefore, it is difficult the draw any conclusion based on the i-band data. The lack of any significant correlation in the sample of ngNLS1s could be either due to a small sample size which is 8 & 6 sources only for W1 and W2 bands (see Fig. 4) or due to large contamination in the variability amplitude by the host galaxies.

### v2

@@ 3.3 Correlation between optical and mid-infrared variability amplitudes
It is expected that IR emission is reprocessed from optical/UV emission (Lu et al., 2016), which means that variability characteristics in the optical bands will be imprinted in the IR wavelength. To investigate such a scenario in the current sample, we selected Seyfert galaxies that exhibit optical variability and have a redshift-corrected variability amplitude in the MIR. For the r-band light curve, the criteria are satisfied by 14 gNLS1s, 8 ngNLS1s, and 9 gBLS1s in both the W1 and W2 bands. However, for the W2 band, only 6 ngNLS1s meet the criteria, while 8 ngNLS1s meet the criteria for the W1 band. Using this sample, we constructed correlation plots between and both and . A representative plot showing the correlated variability amplitudes of the r band optical light curve with the MIR W1 and W2 bands for gNLS1s (left), ngNLS1s (middle), and gBLS1s (right) is presented in Fig. 4. To examine the relationship between MIR and optical variability amplitudes, we employed orthogonal distance regression (ODR) fitting on the data points in the variability amplitude correlation plot, accounting for uncertainties on both axes. Furthermore, we calculated the Pearson rank correlation coefficient () to quantify the correlation between the variability amplitudes for the gNLS1s, ngNLS1s, and gBLS1s samples at both the optical and MIR wavelengths. In this plot, the upper panels illustrate the correlation between and , while the lower panels show the correlation between and . For the r band light curve, the results of the correlation analysis, such as slope, intercept, Pearson r for the W1 and W2 bands, are displayed in the upper left corner of each panel in Fig. 4.
@@
| The number of sources used to estimate the statistics are tabulated inside the bracket ‘[]’. | ||||||||||
@@ 3.4 Color variability
The optical emission observed in AGNs arises from two primary components: the quasi-thermal emission originating from the accretion disk and the nonthermal synchrotron emission from the relativistic jet. Analyzing the color variability of the current sample can provide an effective way to differentiate these components that contribute to the overall flux. Similarly, IR emission, which is considered reprocessed optical/UV radiation (see, Storchi-Bergmann et al., 1992; Lu et al., 2016), offers a unique opportunity to examine the interplay of thermal and nonthermal processes in AGNs. However, to ensure accurate spectral analysis, simultaneous measurements in different bands must be obtained within a 30-minute interval, as described in Sect. 2. This criterion helps mitigate any uncertainties caused by temporal variations and ensures a reliable study of the spectral behavior. Consequently, the color variability analysis in both optical and MIR wavelengths is limited to a subset of sources that meet this stringent timing requirement. The details of the final samples satisfying these conditions are provided at the end of Sect. 2.
@@
For the samples of gNLS1s, ngNLS1s, and gBLS1s mentioned in the last paragraph of Sect. 2, we generated color-magnitude plots, including g-r vs. r, r-i vs. i, and W1-W2 vs. W2, using the available quasi-simultaneous data across different optical and MIR bands. Here, the magnitudes of the longer wavelength are chosen on the X-axis, and the difference between the magnitudes of shorter and longer wavelengths is considered for the color to make the color-magnitude analysis consistent across optical and MIR wavelengths. However, during the color-magnitude analysis, we noticed that its trend reverses when we use shorter wavelength (bluer) magnitudes instead of longer wavelength (redder) on the X-axis, especially for ngNLS1s in the MIR wavelength. To account for this, we produced three combinations of color-magnitude diagrams using m1 and m2 magnitudes, that are, (m1-m2) vs. m2, (m1-m2) vs. m1, and (m1-m2) vs. (m1+m2). Representative examples of these long-term best-fit color-magnitude plots for the gNLS1-J094857.32002225.5, ngNLS1-J144153.98482227.2, and gBLS1-J110249.85525012.7, respectively, are presented in Figs. 6, 7, and 8 corresponding to the color-magnitude combinations (m1-m2) vs. m2, (m1-m2) vs. m1, and (m1-m2) vs. (m1+m2). To examine the relationship between color and magnitude, we employed an ODR fitting on the data points in the color-magnitude plots, accounting for uncertainties in both axes. Furthermore, we calculated to quantify the correlation between color and magnitude for the gNLS1s, ngNLS1s, and gBLS1s samples across both optical and MIR wavelengths.
@@
To qualitatively identify bluer-when-brighter (BWB) or redder-when-brighter (RWB) trends from color-magnitude diagrams, we classified sources based on values, considering indicative of a BWB trend and indicative of an RWB trend. For sources with values of between and 0.5, we assumed that there is no significant correlation (termed NOT) between color and magnitude. The number of sources that exhibit trends of RWB, BWB, or NOT, along with their fractions relative to the sample size, is summarized in Table 3. The slope, intercept, and color trends for the combinations of the color-magnitude diagrams (m1-m2) vs. m2, (m1-m2) vs. m1, and (m1-m2) vs. (m1+m2) of the individual source are tabulated in Table 6.
@@
| gNLS1 [17] | 4 (23.53%) | 5 (29.41%) | 8 (47.06%) | gNLS1 [10] | 5 (50.00%) | 1 (10.00%) | 4 (40.00%) | gNLS1 [22] | 12 (54.54%) | 1 (04.54%) | 9 (40.91%) |
| ngNLS1 [128] | 13 (10.16%) | 31 (24.22%) | 84 (65.62%) | ngNLS1 [74] | 36 (48.65%) | 13 (17.57%) | 25 (33.78%) | ngNLS1 [176] | 49 (27.84%) | 7 (03.98%) | 118 (67.04%) |
| gBLS1 [10] | 8 (80.00%) | 0 (0.00%) | 2 (20.00%) | gBLS1 [6] | 3 (50.00%) | 1 (0.00%) | 2 (50.00%) | gBLS1 [10] | 3 (30.00%) | 3 (10.00%) | 4 (60.00%) |
| The number of sources used to estimate the statistics are tabulated inside the bracket ‘[]’. | |||||||||||
@@ 4 Results and discussion
In this study, we systematically investigated flux and color variability for redshift-matched samples of gNLS1s, ngNLS1s, and gBLS1s in optical and mid-infrared wavelengths. For this purpose, we used high-cadence optical light curves in the g, r, and i bands of ZTF and mid-infrared light curves in the W1 and W2 bands from WISE. To analyze flux variability in the optical wavelength, we initially applied the test using Eq. 1 to the g, r, and i band light curves to determine the presence or absence of intrinsic variability. Subsequently, was calculated exclusively for the variable sources using equation 2, whereas F was calculated for the entire sample, including both variable and non-variable targets, using equation 3. The mean and median values of and are tabulated in Table 1.
@@
From Table 1 and Fig. 2, it is evident that gBLS1s exhibit greater variability, as indicated by both and , compared to gNLS1s across the g, r, and i bands, whether considering mean or median values. This finding is consistent with previous studies on smaller and larger samples of gNLS1s and gBLS1s (Klimek et al., 2004; Ai et al., 2010, 2013; Rakshit & Stalin, 2017). As a class, NLS1s demonstrate lower optical variability amplitudes than BLS1s. This lower variability in NLS1s can be attributed to their higher Eddington ratios (REdd). Since NLS1s accrete at rates significantly higher than BLS1s (Boroson & Green, 1992; Peterson et al., 2000; Ojha et al., 2020), their optical emission is expected to have a relatively stronger thermal component from the accretion disk compared to the synchrotron emission (e.g., see Zhou et al., 2007; Paliya et al., 2014). However, the optical emission from AGNs, dominated by thermal processes, is less variable than the Doppler-boosted synchrotron emission from jets, as a result, the higher contribution of thermal emission in NLS1 due to its higher accretion rate can effectively suppress the amplitude of their optical variability (e.g., see Ojha et al., 2019). However, in contrast to gBLS1s and gNLS1s, ngNLS1s are scarcely variable across the g, r, and i bands with a mean DC () of 5% (see Table 1, see, e.g., Fig. 2), suggesting that their variability is due to instabilities in the accretion disc.
Based on the observed variability characteristics of gBLS1s, gNLS1s, and ngNLS1s, we interpret the differences in their optical variability amplitudes as arising from the relative contributions of jet and accretion disk emission. The gBLS1s exhibit the highest variability, consistent with a dominant jet component and strong Doppler boosting, which enhances both emission and variability. The gNLS1s, while also hosting relativistic jets, likely have a more substantial contribution from the accretion disk due to their high accretion rates, which acts to dilute the observed variability. Consequently, they exhibit lower variability amplitudes than gBLS1s. In contrast, ngNLS1s, which lack jet emission, show the least variability, consistent with a thermal disk-dominated origin. These results suggest that jets enhance optical variability via relativistic beaming, whereas strong thermal emission from the accretion disk, particularly in high-accretion-rate systems such as NLS1s, suppresses it. The weak variability observed in ngNLS1s is therefore likely driven by intrinsic instabilities within the accretion disk. Although it has been proposed that this disk-driven variability may occur on timescales shorter than those probed by our observations, which alone cannot account for the significantly higher variability seen in jet-dominated systems such as gNLS1s and gBLS1s.
| Seyfert | OP | IR | Slope | Intercept | |
| galaxies | band | band | |||
| gNLS1s | g | W1 [14] | 6.092.44 | 0.220.53 | 0.46 |
| W2 [15] | 6.933.00 | 0.290.55 | 0.43 | ||
| r | W1 [14] | 3.780.96 | 0.150.22 | 0.62 | |
| W2 [14] | 3.350.92 | 0.280.19 | 0.50 | ||
| i | W1 [10] | 4.061.70 | 0.060.39 | 0.26 | |
| W2 [10] | 3.301.35 | 0.070.33 | 0.15 | ||
| ngNLS1s | g | W1 [10] | 20.7228.07 | 2.514.30 | 0.02 |
| W2 [6] | 62.76288.79 | 7.8639.26 | 0.23 | ||
| r | W1 [8] | 11.7014.20 | 1.282.26 | 0.07 | |
| W2 [6] | 89.17780.66 | 11.70107.68 | 0.31 | ||
| i | W1 [3] | 2.220.99 | 0.150.16 | 0.89 | |
| W2 [3] | 5.652.81 | 0.240.38 | 0.85 | ||
| gBLS1s | g | W1 [9] | 3.941.19 | 0.070.30 | 0.74 |
| W2 [9] | 4.801.89 | 0.320.48 | 0.68 | ||
| r | W1 [9] | 5.641.61 | 0.330.40 | 0.72 | |
| W2 [9] | 6.772.48 | 0.670.62 | 0.68 | ||
| i | W1 [6] | 1.321.25 | 0.720.50 | 0.18 | |
| W2 [6] | 1.501.26 | 0.680.49 | 0.17 |
| Seyfert | OP | IR | Slope | Intercept | |
| galaxies | band | band | |||
| gNLS1s | g | W1 [14] | 1.010.40 | 0.580.09 | 0.65 |
| W2 [15] | 2.680.80 | 0.120.12 | 0.61 | ||
| r | W1 [14] | 0.810.33 | 0.620.08 | 0.64 | |
| W2 [14] | 0.750.24 | 0.620.06 | 0.50 | ||
| i | W1 [10] | 0.920.35 | 0.260.07 | 0.30 | |
| W2 [10] | 0.790.25 | 0.280.05 | 0.19 | ||
| ngNLS1s | g | W1 [10] | 0.311.45 | 0.720.23 | 0.21 |
| W2 [6] | 0.821.35 | 0.760.20 | 0.03 | ||
| r | W1 [8] | 0.381.70 | 0.480.29 | 0.21 | |
| W2 [6] | 2.622.72 | 0.920.40 | 0.17 | ||
| i | W1 [3] | 5.855.05 | 0.390.74 | 0.87 | |
| W2 [3] | 13.158.64 | 1.171.13 | 0.82 | ||
| gBLS1s | g | W1 [9] | 1.700.54 | 0.350.19 | 0.76 |
| W2 [9] | 1.690.63 | 0.350.22 | 0.71 | ||
| r | W1 [9] | 2.980.69 | 0.060.15 | 0.88 | |
| W2 [9] | 3.270.90 | 0.060.21 | 0.85 | ||
| i | W1 [6] | 0.010.47 | 0.920.24 | 0.27 | |
| W2 [6] | 0.010.45 | 0.910.22 | 0.26 |
| SDSS Name | Galaxy | Slope | Intercept | Color | Plot | |
| type | trends | for | ||||
| 0.240.09 | 4.89 1.86 | 0.53 | RWB | gr vs. r | ||
| 0.000.12 | 0.25 2.41 | 0.10 | NOT | gr vs. g | ||
| 0.070.06 | 2.98 2.24 | 0.34 | RWB | gr vs. gr | ||
| 0.200.20 | 4.26 3.77 | 0.56 | RWB | ri vs. i | ||
| J112758.87362028.4 | gNLS1 | 0.05 0.23 | 0.71 4.50 | 0.24 | NOT | ri vs. r |
| 0.05 0.11 | 2.16 4.37 | 0.42 | RWB | ri vs. ri | ||
| 0.26 0.03 | 4.56 0.48 | 0.53 | RWB | W1W2 vs. W2 | ||
| 0.03 0.04 | 0.55 0.64 | 0.00 | NOT | W1W2 vs. W1 | ||
| 0.06 0.02 | 2.88 0.59 | 0.30 | RWB | W1W2 vs. W1W2 | ||
| 0.75 0.08 | 15.13 1.53 | 0.72 | RWB | gr vs. r | ||
| 0.99 0.22 | 19.90 4.36 | 0.24 | NOT | gr vs. g | ||
| 0.34 0.07 | 13.84 2.86 | 0.37 | RWB | gr vs. gr | ||
| 1.08 0.18 | 21.74 3.64 | 0.79 | RWB | ri vs. i | ||
| J144153.98482227.2 | ngNLS1 | 1.93 0.54 | 38.7910.93 | 0.54 | BWB | ri vs. r |
| 0.71 0.26 | 28.7310.31 | 0.31 | RWB | ri vs. ri | ||
| 1.23 0.13 | 19.21 1.99 | 0.82 | RWB | W1W2 vs. W2 | ||
| 1.77 0.32 | 26.98 5.09 | 0.56 | BWB | W1W2 vs. W1 | ||
| 1.03 0.25 | 32.62 7.62 | 0.35 | RWB | W1W2 vs. W1W2 | ||
| 0.34 0.04 | 6.82 0.86 | 0.67 | RWB | gr vs. r | ||
| J110249.85525012.7 | gBLS1 | 0.29 0.08 | 5.86 1.61 | 0.23 | NOT | gr vs. g |
| 0.17 0.03 | 6.70 1.20 | 0.50 | RWB | gr vs. gr | ||
| 0.26 0.12 | 5.40 2.27 | 0.68 | RWB | ri vs. i | ||
| 0.18 0.18 | 3.87 3.62 | 0.32 | NOT | ri vs. r | ||
| 0.12 0.07 | 4.92 2.93 | 0.54 | RWB | ri vs. ri | ||
| 0.44 0.03 | 7.17 0.49 | 0.64 | RWB | W1W2 vs. W2 | ||
| 0.12 0.05 | 0.98 0.71 | 0.20 | NOT | W1W2 vs. W1 | ||
| 0.11 0.02 | 4.10 0.66 | 0.30 | NOT | W1W2 vs. W1W2 | ||
@@
| Notes: A portion of this table is presented here to display its form and content, however, the entire table | ||||||
| is available in the online electronic version. | ||||||
@@
In contrast to the accretion disc based variability in ngNLS1s, the origin of resulted flux variability of gNLS1s and gBLS1s could either be due to accretion disc instabilities (e.g., Wiita et al., 1991; Chakrabarti & Wiita, 1993; Mangalam & Wiita, 1993) or fluctuations emerging within the jet (e.g., Marscher & Gear, 1985; Wagner & Witzel, 1995; Marscher, 2014). Therefore, we have adopted the variability amplitude cut of 0.4 mags suggested by Bauer et al. (2009) to distinguish between jet-based variability and accretion disc instability-based variability. Thus, we first checked for the variable targets in the samples of gNLS1s and gBLS1s, and found that 15 out of 17 variable gNLS1s (88%) and 10 out of 10 variable gBLS1s (100%) exhibit 0.4 mags for the r-band light-curves. Among the gNLS1s, 89% and 75% of the sample are found to be variable in the g and i bands, respectively. In contrast, all sources (100%) in the gBLS1 sample exhibit variability in both the g and i bands. From this analysis, it appears that gBLS1s tend to be more variable compared to gNLS1s, which could be due to the larger observed contribution of jets more closely aligned towards the observer line of sight in gBLS1s (e.g. Safna et al., 2020). A comparatively small percentage of the variable targets with 0.4 mags in the sample of gNLS1s could also be due to a difference in their accretion rates as discussed above (see 2nd para). On the other hand, a similar kind of variability nature resulted from the samples of gNLS1s, ngNLS1s and gBLS1s in the mid-infrared W1 and W2 bands (see Table 1, Fig. 3), supporting a reprocessing of optical/UV emission from the central engine of Seyfert galaxies by the dusty torus to MIR wavelength. Furthermore, to strengthen the reprocessing scenario found here, we have checked the correlation of with and for optically variable targets having 0.1 in the mid-infrared wavelength for the samples of gNLS1s, ngNLS1s, and gBLS1s.
A strong and statistically significant correlation () is found between the MIR variability amplitude () and both the optical and in the sample of gBLS1s, particularly for the g and r-band light curves with the W1 and W2-band light curves (see Fig. 4; Tables 4, 5). In contrast, while the gNLS1 sample also exhibits statistically significant correlations (), these are consistently weaker than those found for gBLS1s in the same bands. This trend suggests that MIR variability in both gBLS1s and gNLS1s is primarily driven by instabilities in the accretion disk. However, the comparatively weaker correlations in gNLS1s may be due to additional contamination in their optical variability amplitudes, likely arising from stronger host galaxy contributions. Moreover, this interpretation supports the view that MIR variability in gBLS1s is more strongly disk-dominated, whereas in gNLS1s, the combined influence of jets and host contamination may dilute the disk-driven variability signature in the optical bands. Significant correlations are primarily seen in the g and r bands, while correlations involving the i-band light curves are generally weak or statistically insignificant. This could be attributed to the relatively sparse sampling in the i band, making it difficult to draw definitive conclusions from those data. No significant correlations are found for the ngNLS1 sample in any optical band, except for a notable trend in the i band. However, given the small sample size (only 8 and 6 sources for the W1 and W2 bands, respectively), along with potentially strong host galaxy contamination, the lack of significant correlations in ngNLS1s may result from these limitations rather than the absence of an intrinsic physical connection between optical and MIR variability.
Given that the optical , , and bands are sampled more finely than the MIR W1 and W2 bands for our sample of Seyfert galaxies, we investigated whether variability at optical and infrared wavelengths is sensitive to differences in cadence. To do this, we selected sources that were optically variable and had in the MIR, and constructed quasi-simultaneous optical-MIR pairs using windows of days centered on each MIR epoch. Within each window, multiple optical measurements were replaced by their median, with uncertainties propagated via the rms of reported errors. For each , we recomputed the fractional variability amplitude, , in all bands, and quantified the optical-MIR coupling using the newly constructed optical and MIR light curves (e.g., see Fig. 5), applying ODR that accounts for errors on both axes, along with . Increasing modestly raises across all bands, as expected when integrating variability over longer timescales, but does not systematically weaken optical-MIR correlations where they are already present at finer matching. Sources with sparse pairs within d often yield unconstrained results, which become well defined at d, whereas sources with good overlap show stable slopes () and high values () across all windows. We conclude that (i) primarily probes variability on the week–month scale, and (ii) optical and MIR variations occur on comparable timescales. Resampling optical light curves to MIR-like cadence by widening the matching window does not weaken the optical–MIR correlation; if anything, it clarifies it in sparsely sampled cases, indicating that both bands trace the same underlying variability and that the observed cross-band agreement is not a sampling artifact.